# Per-example gradients: a new frontier for understanding and improving optimizers

### v1

@@ Appendix A Just-in-time compilation, tracing, and code optimization
However, not only such an implementation may be too naive, but it is also a naive view of how code is compiled in deep learning frameworks, like JAX (Bradbury et al., 2018) or Pytorch (Paszke et al., 2019). Such frameworks offer just-in-time compilations (jit) techniques that can tailor the compilations to a given backend and optimize the implementation to some extent. Below is the ”jitting” for the average elementwise square.
@@
Just-in-time compilation of a program $\mathcal{P}$ takes typical inputs of $P$, i.e., inputs with the shape and type that $\mathcal{P}$ will be run with, and lists all operations involved in $\mathcal{P}$ in order with the size and types of their inputs and outputs. Such a process is called tracing the program. In JAX, we get then access to the computational graph of $\mathcal{P}$ in the form of a jaxpr that lists operations as JAX primitives in the topological order of the computational graph. Provided with such a computational graph in terms of some primitives, we can define a new computational graph to compute for example gradients of the outputs of $\mathcal{P}$ w.r.t. its inputs. Below is the jaxpr for the average elementwise square.
@@
The computational graph leaves also room for optimizations and tailored implementations for given hardware. For that, jaxprs are first converted in high-level operations (HLO) code, that are intermediate representations. They are independent of the framework (JAX, Pytorch, Tensorflow, etc…), so lower level than JAX. But they are also independent of the hardware, so higher level than targeted code for e.g. GPUs. Below is the hlo code for the average elementwise square.
@@
HLO code can then be parsed to find simplifications or optimizations. Note that finding the high-level optimal implementation of a program may not be possible. The optimization of the implementation is done by parsing the graph with a series of simple rules (some available publicly on the OpenXLA project) that search for specific simplifications. The order in which the simplifications are done may even have an impact on the final implementation (Ganai et al., 2023).
@@ B.1 From broadcasting weights, to reducing gradients, and injecting new statistics
Computing the gradient amounts to composing the Vector-Jacobian Products (VJP) of the operations, i.e., the adjoint [^1] For a linear operator $a:\mathcal{E}\mapsto\mathcal{F}$ from a Euclidean space $\mathcal{E}$ to a Euclidean space $\mathcal{F}$, equipped with respective inner products $\langle\cdot,\cdot\rangle_{\mathcal{E}}$ and $\langle\cdot,\cdot\rangle_{\mathcal{F}}$, the adjoint of $a$ is the unique operator, denoted $a^{*}$ such that for any $u,v\in\mathcal{E}\times\mathcal{F}$, $\langle a(u),v\rangle_{\mathcal{F}}=\langle u,a^{*}(v)\rangle_{\mathcal{E}}$. of their linear approximations on their inputs, in reverse order. Namely, following e.g. Blondel and Roulet (2024, Chapter 2),
@@ B.3 Jaxpr surgery vs vmap
As explained in Appendix A, the native vmap transformation in JAX can benefit from optimizations at the HLO code level or even at a lower level of hardware. So even though the jaxprs of the vmap and the jaxpr surgery in the example above differ, the real performance of these methods can only be judged in the final code execution.
@@
B.4 Memory bottlenecks in deep networks
Section 2.3 summarized memory footprints of gradient computations for dense layers applied to either vector inputs or sequences of vectors. The remark can be summarized and generalized as a simple fact.
Fact B.1.
In a network, for any parameterized operation processing inputs representations $s\in\mathbb{R}^{R}$ using parameters $w\in\mathbb{R}^{P}$, if the dimension of the input is greater than the dimension of the parameters, i.e., $R>P$, then the cost of storing $B$ individual gradients of the loss of the network w.r.t. $w$ is no larger than the checkpointing cost necessary for computing the mini-batch gradients. So if the memory used to store the checkpoints is freed to store the $B$ individual gradients, the peak memory usage of computing the mini-batch gradients is the same as the memory cost to compute the individual gradients and then averaging them.
Proof.
Computing the mini-batch gradient requires checkpointing $B$ inputs $s$ at a cost $BD$. Storing $B$ gradients take $BP$ memory. Memory allocations along program execution are well explained by e.g. Griewank and Walther (2008, Chapter 4). ∎
Fact B.1 may still lower bound the memory necessary for computing the mini-batch loss. Some non-parameterized operations may require extensive additional memory. This is typically the case for attention blocks in transformers that require a priori $BL^{2}$ additional memory to checkpoint the inputs of the activation module.
We can parse common layers through this fact.
  1. 1.
    For dense layers operating on vectors, i.e., of the form $t^{\top}=s^{\top}W$, with $t\in\mathbb{R}^{D},s\in\mathbb{R}^{D}$, $W\in\mathbb{R}^{D\times D}$, Fact B.1 does not apply as we have $R=D<D^{2}=P$.
  2. 2.
    For dense layers operating on sequences of vectors, i.e., of the form $T=SW$, for $T\in\mathbb{R}^{L\times D},S\in\mathbb{R}^{L\times D}$, $W\in\mathbb{R}^{D\times D}$, as long as $L$ the length of the sequence is less than $D$, the width, Fact B.1 applies.
    • This may apply for example in the dense layers of the multihead attention and the mlp blocks of a transformer.
    • This would not apply to the embedding layer and the final head of a transformer whose dimensions depend on the vocabulary that generally has a very large size.
  3. 3.
    Convolutional layers operating on images of size $HWC$ for $H,W,C$ the height, width and number of channels of the image. The parameters are filters of size $K^{2}C^{2}$, with $K$ the kernel size. If $HW>KC$, Fact B.1 applies.
The first point of Fact B.1 is easy to check. The second point depends on the compilation of the program and compiler optimizations (Appendix A). Fact B.1 simply hints that computations involving individual gradients may not be prohibitive for some architectures given an efficient just-in-time compiler.
Figure 2 shows the memory allocated along program execution for the Adam and MicroAdam algorithms on a transformer. We observe that the peak memory is the same. It corresponds well to the peak activation required for reverse-mode automatic differentiation (compare to Figure 8.4 in (Blondel and Roulet, 2024)).
The unchanged peak memory usage may be surprising as individual gradients of the embedding layer may still account for an important part of the computational cost. However, in any case, just-in-time compilation in JAX (Appendix A) may automatically trade-off memory and computational usages.
B.5 Mini-batch gradient structure
@@
For convolutions, several implementations pass by folding/unfolding the images as explained by Bu et al. (2022). Essentially, for an input image $s_{i}\in\mathbb{R}^{H\times W\times C}$ convolved with a filter $w\in\mathbb{R}^{K\times K\times C\times C}$, the image is unfolded into $\sigma_{i}\in\mathbb{R}^{HW\times CK^{2}}$, and the weights into $\omega\in\mathbb{R}^{CK^{2}\times C}$ such that the matrix product $\tau_{i}=\sigma_{i}\omega\in\mathbb{R}^{HW\times C}$ corresponds to the convolution at each output index (variable heights, width, channels, and kernel sizes can be treated with the same logic, see Bu et al. (2022)). The output image is obtained by folding $\tau_{i}$ into $t_{i}\in\mathbb{R}^{H\times W\times C}$. For a batch of examples, the core operation remains a tensor contraction
@@
that can be written as a sum of rank-one vectors. As explained in 2.2 we can exploit this structure for factorable operations like the elementwise square.
@@
For some operations weights are preprocessed before being applied to the inputs. For example, weight matrices may be orthonormalized before being applied (Mhammedi et al., 2017). Operations on the inputs take then the form
@@ D.1 Argument for the square root scaling rule
The square root scaling rule can be made more formal in various stochastic differential equation (SDE) limits of training dynamics. Currently such limits have been formally derived for SignSGD and not Adam, but the same square root learning rule applies in both cases. In such a limit one can derive an SDE in continuous time whose dynamics can be mapped onto the discrete stochastic dynamics such that averages of observables match between the two descriptions, with error shrinking with some small parameter (learning rate in the approach of Compagnoni et al. (2025), inverse dimension in Xiao et al. (2025)).
In our training setup, we can see that the square root learning rule induces similar learning curves in terms of number of samples over some range of $B$ (Figure 9). We first found the best learning rate $\eta^{*}_{B_{0}}$ for $B_{0}=64$. We then use the square root scaling rule to generate learning rates for other batch sizes with $0.5\eta^{*}_{B_{0}}$ (Figure 9, left), and $\eta^{*}_{B_{0}}$ (Figure 9, right). We see that there is a crossover from small $B=8$ to the larger batch sizes where the learning curve is quite different. From $B=16$ to $B=128$ the intermediate to late time learning curves are similar, while for larger batch sizes (and therefore learning rates) the curves become non-universal and eventually diverge. Smaller learning rates show better agreement as predicted by the theory; the heuristic (and the SDE equivalents) break down if the gradient distribution changes over a few number of steps. This can happen because of effects like feature learning, or even more simply by making significant progress towards the objective.
@@ D.3 Adam $\mu^{2}$ and $\sigma^{2}$ statistics
In order to further understand the consistency of the square root learning rule with our observation that $\mu^{2}$ seems to dominate the preconditioner at early times, we measured $\mu^{2}$ and $\sigma^{2}$ using the estimators from 8 for models trained with varying batch size. Plotting the ratio of $\hat{\mu}^{2}$ to $\hat{\sigma}_{eff}^{2}$ ($\sigma^{2}/B$ for regular Adam, $\sigma^{2}$ for MicroAdam) we see that at all batch sizes we studied, $\mu^{2}$ dominates at early times for Adam (Figure 7, left). The ratios and their dynamics were somewhat consistent even for Adam, where $\hat{\sigma}^{2}_{eff}$ has explicit dependence on $B$.
@@ Appendix E Experimental details
We train a standard decoder-only transformer from the Nanodo codebase (Liu et al., 2024).
E.1 Dataset and tokenizer
We train on the C4 dataset (Raffel et al., 2020), using use the SentencePiece tokenizer from Raffel et al. (2020).
E.2 Architecture
The architecture consists in $N$ identical transformer blocks. Each block consists in a multi-head causal dot-product self-attention layer followed by an MLP layer. Layer normalization is applied to both the input of the attention and the input of the MLP layer. The MLP layer uses a gelu activation function. The model employs a learned positional embedding. The model uses an independent layer to map back to the vocabulary space. So in plain code, the architecture reads
@@ E.2 Architecture
E.3 Optimization generic details
Below, we present generic hyperparameters’ settings. Specific setups are detailed in
@@ E.3 Optimization generic details
In all experiments, the number of steps is fixed as (following Nanodo’s recipe (Liu et al., 2024) which follows Chinchilla’s (Hoffmann et al., 2022) scaling factor $c=20$)
@@
E.4 Experimental specific details.
@@
For the scaling experiments (Figure 4, 4), as the batch size varies, the total number of steps vary according to the rule $K=\lfloor cP/T\rfloor$ with $T=LB$. The number of warmup steps is multiplied by $B/1024$ as it had been selected for batch size of 1024. As the base learning rate was chosen for a batch size $B=1024$, the learning rate is adjusted as $\eta_{\gamma}=\gamma\eta$ with a factor $\gamma=B/1024$ for all variants except Adam that uses $\gamma=\sqrt{B/1024}$. This same factor is used to adjust the weight decay as $\omega_{B}=\omega_{\gamma}$.
@@
We implemented MicroAdamVar by using two EMA estimators $\nu_{\rm adam}$ and $\nu_{\rm micro}$, and combining them in accordance with Equation 8. We then carried out a similar bath size scaling experiment to Figure 4.
@@ Appendix F Detailed pseudocode
Algorithm 1 MicroAdam algorithms
@@
6: Preconditioner computation MicroAdamX
@@
12:  Compute average gradient $\bar{g}_{t}=\frac{1}{B}\sum_{i=1}^{B}\nabla f(\theta_{t-1};x_{i}^{(t)})$
13:  Update EMA of gradients $\mu_{t}=\beta_{1}\mu_{t-1}+(1-\beta_{1})\bar{g}_{t}$
14:  if MicroAdamX = MicroAdam then
15:   Compute preconditioner as average element-wise square  $\nu=\frac{1}{B}\sum_{i=1}^{B}g_{i}^{2}$ for $g_{i}=\nabla f(\theta_{k};x_{i}^{(t)})$
16:  else if MicroAdamX = MicroAdamVar then
17:   Compute preconditioner as element-wise estimate variance  $\nu=\frac{B}{1-B}\left(\frac{1}{B}\sum_{i=1}^{B}g_{i}^{2}-\left(\frac{1}{B}\sum_{i=1}^{B}g_{i}\right)^{2}\right)$ for $g_{i}=\nabla f(\theta_{k};x_{i}^{(t)})$
18:  else if MicroAdamX = MicroAdamMSQ then
19:   Compute preconditioner as element-wise estimate variance  $\nu=\frac{B}{1-B}\left(\left(\frac{1}{B}\sum_{i=1}^{B}g_{i}\right)^{2}-\frac{1}{B}\sum_{i=1}^{B}g_{i}^{2}\right)$ for $g_{i}=\nabla f(\theta_{k};x_{i}^{(t)})$
20:  end if
21:  Update EMA of preconditioner $\nu_{t}=\beta_{2}\nu_{t-1}+(1-\beta_{2})\nu$
22:  Apply bias corrections on both EMAs: $\hat{\mu}_{t}=\mu_{t}/(1-\beta_{1}^{t})$,  $\hat{\nu}_{t}=\nu_{t}/(1-\beta_{1}^{t})$
23:  if MicroAdamX = MicroAdamMSQ then
24:   $\hat{\nu}_{t}\leftarrow\max(0,\hat{\nu}_{t})$
25:  end if
26:  Define update direction $u_{t}=\hat{\mu}_{t}/\sqrt{\varepsilon+\hat{\nu}_{t}}$
27:  Apply update $\theta_{t}=\theta_{t-1}-\eta_{t}u_{t}$
28:end for

### v2

@@ Appendix A Just-in-time compilation, tracing, and code optimization
However, not only such an implementation may be too naive, but it is also a naive view of how code is compiled in deep learning frameworks, like JAX (jax2018) or Pytorch (pytorch2019). Such frameworks offer just-in-time compilations (jit) techniques that can tailor the compilations to a given backend and optimize the implementation to some extent. Below is the ”jitting” for the average elementwise square.
@@
Just-in-time compilation of a program $\mathcal{P}$ takes typical inputs of $P$, i.e., inputs with the shape and type that $\mathcal{P}$ will be run with, and lists all operations involved in $\mathcal{P}$ in order with the size and types of their inputs and outputs. Such a process is called tracing the program. In JAX, we get then access to the computational graph of $\mathcal{P}$ in the form of of a jaxpr that lists operations as JAX primitives in the topological order of the computational graph. Provided with such a computational graph in terms of some primitives, we can define a new computational graph to compute for example gradients of the outputs of $\mathcal{P}$ w.r.t. its inputs. Below is the jaxpr for the average elementwise square.
@@
The computational graph leaves also room for optimizations and tailored implementations for given hardwares. For that, jaxprs are first converted in high-level operations (HLO) code, that are intermediate representations. They are independent of the framework (JAX, Pytorch, Tensorflow, etc…), so lower level than JAX. But they are also independent of the hardware, so higher level than targeted code for e.g. GPUs. Below is the hlo code for the average elementwise square.
@@
HLO code can then be parsed to find simplifications or optimizations. Note that finding the high-level optimal implementation of a program may not be possible. The optimization of the implementation is done by parsing the graph with a series of simple rules (some available publicly on the OpenXLA project) that search for specific simplifications. The order in which the simplifications are done may even have an impact on the final implementation (ganai2023target).
@@ B.1 From broadcasting weights, to reducing gradients, and injecting new statistics
Computing the gradient amounts to composing the Vector-Jacobian Products (VJP) of the operations, i.e., the adjoint [^1] For a linear operator $a:\mathcal{E}\mapsto\mathcal{F}$ from a Euclidean space $\mathcal{E}$ to an Euclidean space $\mathcal{F}$, equipped with respective inner products $\langle\cdot,\cdot\rangle_{\mathcal{E}}$ and $\langle\cdot,\cdot\rangle_{\mathcal{F}}$, the adjoint of $a$ is the unique operator, denoted $a^{*}$ such that for any $u,v\in\mathcal{E}\times\mathcal{F}$, $\langle a(u),v\rangle_{\mathcal{F}}=\langle u,a^{*}(v)\rangle_{\mathcal{E}}$. of their linear approximations on their inputs, in reverse order. Namely, following e.g. blondel2024elements,
@@ B.3 Jaxpr surgery vs vmap
As explained in Appendix A, the native vmap transformation in JAX can benefit from optimizations at the HLO code level or even at a lower level of hardware. So even though the jaxprs of the vmap and the jaxpr surgery in the example above differ, the real performance of these methdos can only be judged in the final code execution.
@@
B.4 Mini-batch gradient structure
@@
For convolutions, several implementations pass by folding/unfolding the images as explained by bu2022scalable. Essentially, for an input image $s_{i}\in\mathbb{R}^{H\times W\times C}$ convolved with a filter $w\in\mathbb{R}^{K\times K\times C\times C}$, the image is unfolded into $\sigma_{i}\in\mathbb{R}^{HW\times CK^{2}}$, and the weights into $\omega\in\mathbb{R}^{CK^{2}\times C}$ such that the matrix product $\tau_{i}=\sigma_{i}\omega\in\mathbb{R}^{HW\times C}$ corresponds to the convolution at each output index (variable heights, width, channels, and kernel sizes can be treated with the same logic, see bu2022scalable). The output image is obtained by folding $\tau_{i}$ into $t_{i}\in\mathbb{R}^{H\times W\times C}$. For a batch of examples, the core operation remains a tensor contraction
@@
that can be written as a sum of rank-one vectors. As explained in 2.3 we can exploit this structure for factorable operations like the elementwise square.
@@
For some operations weights are preprocessed before being applied to the inputs. For example, weight matrices may be orthonormalized before being applied (mhammedi2017efficient). Operations on the inputs take then the form
@@ D.1 Argument for the square root scaling rule
The square root scaling rule can be made more formal in various stochastic differential equation (SDE) limits of training dynamics. Currently such limits have been formally derived for SignSGD and not Adam, but the same square root learning rule applies in both cases. In such a limit one can derive an SDE in continuous time whose dynamics can be mapped onto the discrete stochastic dynamics such that averages of observables match between the two descriptions, with error shrinking with some small parameter (learning rate in the approach of compagnoni2024adaptive, inverse dimension in xiao2025exact).
In our training setup, we can see that the square root learning rule induces similar learning curves in terms of number of samples over some range of $B$ (Figure 10). We first found the best learning rate $\eta^{*}_{B_{0}}$ for $B_{0}=64$. We then use the square root scaling rule to generate learning rates for other batch sizes with $0.5\eta^{*}_{B_{0}}$ (Figure 10, left), and $\eta^{*}_{B_{0}}$ (Figure 10, right). We see that there is a crossover from small $B=8$ to the larger batch sizes where the learning curve is quite different. From $B=16$ to $B=128$ the intermediate to late time learning curves are similar, while for larger batch sizes (and therefore learning rates) the curves become non-universal and eventually diverge. Smaller learning rates show better agreement as predicted by the theory; the heuristic (and the SDE equivalents) break down if the gradient distribution changes over a small number of steps. This can happen because of effects like feature learning, or even more simply by making significant progress towards the objective.
@@ D.3 Adam $\mu^{2}$ and $\sigma^{2}$ statistics
In order to further understand the consistency of the square root learning rule with our observation that $\mu^{2}$ seems to dominate the preconditioner at early times, we measured $\mu^{2}$ and $\sigma^{2}$ using the estimators from 3.2.2 for models trained with varying batch size. Plotting the ratio of $\hat{\mu}^{2}$ to $\hat{\sigma}_{\text{eff}}^{2}$ ($\sigma^{2}/B$ for regular Adam, $\sigma^{2}$ for MicroAdam) we see that at all batch sizes we studied, $\mu^{2}$ dominates at early times for Adam (Figure 8, left). The ratios and their dynamics were somewhat consistent even for Adam, where $\hat{\sigma}^{2}_{\text{eff}}$ has explicit dependence on $B$.
@@ Appendix E Experimental details
E.1 Testing per-example methods across workloads
For the experiments on the Algoperf codebase, we implemented MicroAdam in JAX using vmap via the Algoperf API. We trained on $2\times 4$ slices of TPU v5e. We removed batchnorm from any workloads which had it since this normalization is incompatible with a vmap based approach. All workloads were run with the competition configurations. For convenience, the batch sizes (raw and per device) for each workload group can be found in Table 1.
Workload name Batch size Batch size per device
fastmri unet 32 4
fineweb lm 64 8
imagenet (all) 1024 128
wmt transformer 128 16
librispeech (all) 256 32
Table 1: Algoperf workload batch sizes.
E.2 Transformer-specific experiments
In this subsection, we comment on the experiments detailed in Section 3.
E.2.1 Dataset and Tokenizer
For the experiments in Section 3, we train on the C4 dataset (raffel2020exploring), using use the SentencePiece tokenizer from raffel2020exploring.
E.2.2 Architecture
We trained a standard decoder-only transformer from the Nanodo codebase (nanodo), using TPU v5e. The architecture consists in $N$ identical transformer blocks. Each block consists in a multi-head causal dot-product self-attention layer followed by a MLP layer. Layer normalization is applied to both the input of the attention and the input of the MLP layer. The MLP layer uses a gelu activation function. The model employs a learned positional embedding. The model uses an independent layer to map back to the vocabulary space. So in plain code, the architecture reads
@@ E.2.2 Architecture
E.2.3 Optimization generic details
Below, we present generic hyperparameters’ settings. Specific setups are detailed in E.2.4.
@@ E.2.3 Optimization generic details
In all experiments, the number of steps is fixed as (following Nanodo’s recipe (nanodo) which follows Chinchilla’s (hoffmann2022training) scaling factor $c=20$)
@@
E.2.4 Experimental specific details.
@@
For the scaling experiments (Figure 5, 5), as the batch size varies, the total number of steps vary according to the rule $K=\lfloor cP/T\rfloor$ with $T=LB$. The number of warmup steps is multiplied by $B/1024$ as it had been selected for batch size of 1024. As the base learning rate was chosen for a batch size $B=1024$, the learning rate is adjusted as $\eta_{\gamma}=\gamma\eta$ with a factor $\gamma=B/1024$ for all variants except Adam that uses $\gamma=\sqrt{B/1024}$. This same factor is used to adjust the weight decay as $\omega_{B}=\omega_{\gamma}$.
@@
We implemented MicroAdamVar by using two EMA estimators $\nu_{\rm adam}$ and $\nu_{\rm micro}$, and combining them in accordance with Equation 3.2.2. We then carried out a similar batch size scaling experiment to Figure 5.
@@ Appendix F Detailed pseudocode
Algorithm 1 Various Adam algorithms
@@
6:  Adam variant AdamVariant
@@
12:  Compute average gradient
13:   $\mu_{t}=\frac{1}{B}\sum_{i=1}^{B}g_{i,t}$ for $g_{i,t}=\nabla f(\theta_{t};x_{i}^{(t)})$
14:  Update EMA of gradients
15:   $m_{t}=\beta_{1}m_{t-1}+(1-\beta_{1})\mu_{t}$
16:  if AdamVariant is Original then
17:   Compute preconditioner as square of average gradients
18:    $\nu_{t}=\mu_{t}^{2}=\left(\frac{1}{B}\sum_{i=1}^{B}g_{i,t}\right)^{2}$
19:  else if AdamVariant is MicroAdam then
20:   Compute preconditioner as average element-wise square
21:    $\nu_{t}=\frac{1}{B}\sum_{i=1}^{B}g_{i,t}^{2}$
22:  else if AdamVariant is MicroAdamVar then
23:   Compute preconditioner as element-wise estimate variance
24:    $\nu_{t}=\frac{B}{1-B}\left(\frac{1}{B}\sum_{i=1}^{B}g_{i,t}^{2}-\left(\frac{1}{B}\sum_{i=1}^{B}g_{i,t}\right)^{2}\right)$
25:  else if AdamVariant is MicroAdamMSQ then
26:   Compute preconditioner as element-wise estimate squared expectation
27:    $\nu_{t}=\frac{B}{1-B}\left(\left(\frac{1}{B}\sum_{i=1}^{B}g_{i,t}\right)^{2}-\frac{1}{B}\sum_{i=1}^{B}g_{i,t}^{2}\right)$
28:  end if
29:  Update EMA of preconditioner
30:   $v_{t}=\beta_{2}v_{t-1}+(1-\beta_{2})\nu_{t}$
31:  Apply bias corrections on both EMAs:
32:   $\hat{m}_{t}=m_{t}/(1-\beta_{1}^{t})$,
33:   $\hat{v}_{t}=v_{t}/(1-\beta_{2}^{t})$
34:  if MicroAdamX = MicroAdamMSQ then
35:   $\hat{v}_{t}\leftarrow\max(0,\hat{v}_{t})$
36:  end if
37:  Define update direction $u_{t}=\hat{m}_{t}/\sqrt{\varepsilon+\hat{v}_{t}}$
38:  Apply update $\theta_{t}=\theta_{t-1}-\eta_{t}u_{t}$
39:end for
Authors contributions
  • Atish came up with the initial idea of the study, did the theoretical analysis, did some experiments, and contributed to the code.
  • Vincent came up with the computational approach, wrote the majority of the codebase and conducted the majority of the experiments.
  • Atish and Vincent contributed to the scientific questions and wrote the paper.